### v2

# GraphRAFT: Retrieval Augmented Fine-Tuning for Knowledge Graphs on Graph Databases

Abstract
Large language models have shown remarkable language processing and reasoning ability but are prone to hallucinate when asked about private data. Retrieval-augmented generation (RAG) retrieves relevant data that fit into an LLM’s context window and prompts the LLM for an answer. GraphRAG extends this approach to structured Knowledge Graphs (KGs) and questions regarding entities multiple hops away. The majority of recent GraphRAG methods either overlook the retrieval step or have ad hoc retrieval processes that are abstract or inefficient. This prevents them from being adopted when the KGs are stored in graph databases supporting graph query languages. In this work, we present GraphRAFT, a retrieve-and-reason framework that finetunes LLMs to generate provably correct Cypher queries to retrieve high-quality subgraph contexts and produce accurate answers. Our method is the first such solution that can be taken off-the-shelf and used on KGs stored in native graph DBs. Benchmarks suggest that our method is sample-efficient and scales with the availability of training data. Our method achieves significantly better results than all state-of-the-art methods across all four standard metrics on two set of challenging Q&As on large text-attributed KGs.
1 Introduction
Large language models (LLMs) have shown remarkable ability to reason over natural languages. However, when prompted with questions regarding new or private knowledge, LLMs are prone to hallucinations[huang2025survey]. A popular approach to remediate this issue is Retrieval-Augmented Generation (RAG) [lewis2020retrieval]. Given a natural language question and a set of text documents, RAG retrieves a small set of relevant documents and feeds them into an LLM.
In addition to unstructured data such as text documents, many of the real-world structured data come in the form of Knowledge Graphs (KGs). KGs can be used to model a variety of domains such as the Web, social networks, and financial systems. Knowledge Base Question Answering (KBQA)[lan2021survey] studies the methods used to answer questions about KGs. Most of the work studies multi-hop questions on the Web (open-domain)[Yih2016TheVO, freebase, talmor-berant-2018-web].
Historically, there are two categories of approaches for KBQA. One approach is based on semantic parsing[yih-etal-2015-semantic, luo-etal-2018-knowledge, 10.1007/978-3-319-93417-4_46, 10.1145/2806416.2806472, chen-etal-2019-uhop, 10.1145/3357384.3358033, lan-jiang-2020-query], which involves converting natural language questions into logical forms that are often SPARQL[sparql]. Another approach is based on information retrieval[sun2019pullnetopendomainquestion, sun-etal-2018-open, xiong-etal-2019-improving, he-etal-2017-generating], which involves embedding and retrieving triplets from KGs and reranking the extracted data.
There is a growing demand to leverage the language understanding and reasoning power of LLMs to improve open-domain KBQA. At the same time, GraphRAG extends the RAG setup to private KGs. The methodologies began to merge, despite the different application areas. A unifying architecture follows the general framework of 1) identifying some entities in the given question, 2) retrieving a subgraph and 3) prompting an LLM for reasoning.
[baek2023knowledgeaugmentedlanguagemodelprompting, hu2024graggraphretrievalaugmentedgeneration] retrieves and reranks the top nodes (or k-hop ego-subgraphs) and prompts the LLM. [luo2024reasoning] prompts an LLM to generate relation path templates and retrieves concrete paths using in-memory breadth-first search and fintunes the LLM. [sun2024thinkongraph, ma2024think, sui2024fidelisfaithfulreasoninglarge, wang2025reasoning, xia2024knowledgeawarequeryexpansionlarge] identify starting entities and iteratively perform the retrieval step-by-step by prompting the LLM what is the next step to take. [he2024gretriever] prunes a given subgraph with parameterised graph algorithms and prompts an LLM with a textualised version of the local subgraph. [wang2023knowledgedrivencotexploringfaithful] extends LLM chain-of-thought (CoT) by iteratively retrieving from the KG. [chen2025knowledge] enhances the quality of the QA training data.
Although significant progress has been made, there are several limitations of existing methods. Many of the real-world KGs are stored in native graph DBs. Graph DBs come with query engines that optimise user queries, such as Cypher. However, none of the existing methods leverage such graph DBs at all. Any method that explicitly requires step-by-step retrieval (due to iteratively prompting the LLMs) prevents direct multi-hop querying on DBs. This implies that each step of the retrieval process needs to be fully materialised and the fixed sequential order of traversal implementation is frequently suboptimal. Benchmarking is often performed with in-memory graphs represented as tensors with adhoc implementations of subgraph retrieval.
On the other hand, many methods that involve multi-hop retrieval as a single step remain abstract. They usually stop at the point of specifying some path patterns. Converting the abstract path definitions into queries remains a separate challenging task, irrespective of specific query languages[lei2025spider20evaluatinglanguage, texttosparql, brei2024leveragingsmalllanguagemodels, ozsoy2024text2cypherbridgingnaturallanguage]. Another stream of work explicitly requires the questions to be of some logic fragment and converts them into SPARQL. There is also no general way of guaranteeing that the generated queries are syntactically and semantically correct with respect to specific KGs.
In this paper, we propose a simple, modular, and reliable approach for question answering on private KGs (GraphRAG). Specifically, the contributions of our papers are:
  • We propose a new method of fine-tuning an LLM to generate optimal Cypher queries that only requires textual QAs as training data but not user-crafted ground-truth queries.
  • At inference time, we deploy a novel grounded constrained decoding strategy to generate provably syntactically and semantically correct Cypher queries.
  • Our method follows a modular and extensible retrieve-and-reason framework and can be taken off-the-shelf to perform GraphRAG on native Graph DBs.
In addition, benchmarks show that our method achieves significantly beyond SOTA results across all 4 metrics on STaRK-prime and STaRK-mag, two large text-attributed KGs with Q&As for GraphRAG. For example, our method achieves 63.71% Hit@1 and 68.99% Mean Reciprocal Rank (MRR), which is an improvement of 22.81 and 17.79 percentage points, respectively, over the best previous baselines on STaRK-prime. Our method is sample-efficient, achieving beyond SOTA results on all metrics on STaRK-prime when trained with only 10% of the data, and scales with more training data.
2 Related Work
RAG. Recent work extends RAG[lewis2020retrieval] to broader settings. RAFT[raft2024] studies domain-specific retrieval-augmented fine-tuning on by sampling relevant and irrelavant documents. GraphRAG[edge2024local] generalises RAG to global text summarisation tasks in multiple documents. The term GraphRAG has also been widely adopted to mean RAG on (knowledge) graphs, which is the problem setup we study in this paper. [graft] applies RAFT for GraphRAG on document retrieval and summarisation. There are many methodological and implementation improvements[li2024simple] and industrial applications of GraphRAG[Xu_2024].
Message Passing Graph Neural Networks and Graph Transformers. Message-passing Graph Neural Networks (GNNs) iteratively aggregates and updates embeddings on nodes and edges using the underlying graph as the computation graph, hence incorporating strong inductive bias. GNNs have been used within the framework of GraphRAG to improve retrieval[mavromatis2024gnn] or to enhance the graph processing power of LLMs[zhang2022greaselm, yasunaga2021qa]. Alternative model architectures such as Graph Transformers, which employ graph-specific tokenization or positional embedding techniques, have been developed[ying2021transformersreallyperformbad, tokengt]. There are also work[zhao2023learning, he2024harnessing, jiang2024ragraph] that leverage LLMs to improve classical graph problems such as node classification and link prediction on text-attributed graphs.
3 Preliminary
3.1 Graph database and graph query language
Graph DB treats both nodes and edges as primary citizens of the data model. A common graph data model is Labeled Property Graph (LPG). LPGs support types and properties on nodes and edges in a flexible way. For example, nodes of the same type are allowed to have different properties, and nodes and edges can share the same types. LPGs can be used to model virtually all real-world KGs.
A Graph DB stores LPGs efficiently on disks. It also comes with a query engine that allows one to query the graph. A widely used query language is Cypher[cypher][cyphersemantics] (and a variant openCypher[opencypher]), which are implementations of the recent GQL[gql][gqlpaper] standard. We will refer to the two interchangeably throughout this paper. The key ingredient of Cypher is graph pattern matching. A user can query the graph by matching on patterns (e.g paths of a certain length) and predicates (e.g filtering on node and edge properties). An example Cypher query is provided in Figure 1. The execution of a query is carried out by the query engine which heavily optimises the order of executing low-level operators.
An alternative graph data model is Resource Description Framework (RDF)[rdf]. It was originally created to model the Web. A graph is modeled as a collection of triples, commonly referred to as subject-predicate-object, where each element can be identified with a URI (mimicing the Web). Query languages on top of RDFs include SPARQL[sparql]. RDFs also support ontology languages such as Web Ontology Language (OWL)[owl] to model and derive facts through formal logic. Much of the KBQA literature has taken inspiration from RDFs by defining a graph as a collection of triples and reasoning over it in a formal style. Widely used benchmarks such as WebQSP[Yih2016TheVO] are exactly questions about the Web that are answerable by SPARQL.
Many open-source and commercial graph DBs support Cypher over LPGs, such as Neo4j[neo4j], ArangoDB[arangoDB], TigerGraph[tigergraph]. There are also popular RDF stores such as Neptune[neptune], which also supports Cypher over RDFs.
UNWIND \$source_names AS source_name
MATCH (src {name: source_name})-[r1]-(var)-[r2]-(tgt) WHERE src <> tgt
RETURN labels(src)[1] AS label1, src.name AS name1, type(r1) AS type1,
labels(var)[1] AS label2, type(r2) AS type2, labels(tgt)[1] AS label3,
size([t IN collect(DISTINCT tgt) WHERE t.nodeId in \$target_ids | t]) AS
correctCnt, count(DISTINCT tgt) AS totalCnt
Figure 1: An example Cypher query. It takes as user input a list variable source_names and another list target_ids. It iterates through them and find all two-hop neighbours of each source_name node, requiring the second-hop node to be distinct to the source. The query returns aggregate information of the subgraph such as labels and types of nodes and edges, and arithmetic over how many second-hop nodes have ids that are in the user-defined node id list.
3.2 LLM
Modern LLMs are usually auto-regressive decoder-only Transformers as backbones that are trained on the Web. An LLM $f_{\theta}$ has a fixed vocabulary set $\mathcal{V}$. Given a sequence of characters $c_{0},\cdots,c_{n}$, a tokenizer converts it to a sequence of tokens $t_{0},\cdots,t_{k}$ where $t_{i}\in\mathbb{R}^{d}$ and $d$ is the embedding dimension. The tokenizer often compresses multiple characters into one token and splits a word into multiple tokens. Given $t_{0},\cdots,t_{k}$, suppose $f_{\theta}$ has generated tokens $t_{k+1},\cdots,t_{k+n}$, it computes the logits for the $k+n+1^{\text{th}}$ token as $l_{k+n+1}=f_{\theta}(t_{0};t_{k+n})$ where $l_{k+n+1}\in\mathbb{R}^{|\mathcal{V}|}$. The probability of generating any token $x$ is obtained by applying softmax to the logits, $p(t_{k+n+1}=x|t_{0};t_{k+n})=exp(l_{k+n+1}^{x})/\sum_{y\in\mathcal{V}}{exp(l_{k+n+1}^{y})}$. Greedy decoding picks the token with the highest probability at each step. Beam search with width $m$ keeps $m$ sequences of tokens with the highest product of probabilities so far. The generative process normally terminates when some end-of-sequence <eos><eos> token is decoded. Pretrained LLMs can be finetuned efficiently using techniques such as LoRA[hu2022lora], optimising the product of conditional probabilities of next-token prediction.
4 Approach
Let $\mathcal{G}=(V,E,L,l_{v},l_{e},K,W,p_{v},p_{e})$ be a KG stored in a graph DB. $V,E$ are nodes and edges. $L$ is a set of labels. $l_{v}:V\rightarrow\mathcal{P}(L)$ the label assignment function for $V$. $K,W$ the set of property keys and values. $p_{v}:V\rightarrow K\times W$ the property key-value assignment function on the nodes. $l_{e},p_{e}$ are defined analogously for edges. Note that we support the most flexible definition of KGs, where nodes and edges can share labels, and those of the same label can have different properties keys. We assume that each node has at least one property that is text, which is common in the GraphRAG setup. All nodes $v$ are equipped with an additional property which we abbreviate as $z_{v}$, which is the text embedding produced by some text embedder $LM$ on the text attribute.
Given a set of training QAs $\{Q_{i},A_{i}\}$ where $A_{i}\subset V$, we want our model to produce good answers $A_{j}$ for unseen questions $Q_{j}$} according to a variety of metrics. Our approach consists of several steps. First, we create a set of training question-Cypher pairs $\{Q_{i},C_{i}\}$ to finetune a LLM (Section 4.1). At inference time, we deploy a simple method of constrained decoding that is grounded to $\mathcal{G}$ to guarantee syntactically and semantically optimal Cypher queries (Section 4.2). The optimal Cypher is executed to retrieve a text-attributed subgraph for each question and a second LLM is finetuned to jointly reason over text and subgraph topology to select the final answers (Section 4.3).
4.1 Synthesize ground-truth cypher
For a given $Q_{i}$, we few-shot prompt an LLM $L_{0}$ to identify the relevant entities (strings) $n_{1},\cdots n_{k}$ mentioned in the question. In order to address the inherent linguistic ambiguity and noise present in the questions and the graph, if a generated node name does not correspond to an existing node, we perform an efficient vector similarity search directly using vector index at the database level to retrieve the most similar nodes.
$v_{i}=cossim(LM(n_{i}),\{z_{v}\}_{v\in V})$ (1)
This enables more accurate identification of entities than performing native vector similarity between all nodes in the graph and an embedded vector representing the question.
We then construct several Cypher templates that match multi-hop subgraphs around the identified entities $v_{1},\cdots,v_{k}$ with filters on node and relationship types. For each Cypher query $C_{j}$ we execute the query and compute the hits in the answers $A_{i}$}, as well as the number of total nodes. The calculation (as an aggregation step) is performed as part of the query itself which means the nodes and edges and their properties do not need to be materialised and retrieved and hence saves memory and I/O workload. An example such query is shown in Figure 1.
We then rank the queries according to their hits and number of results and retain the top query $C_{i}$. This gives us a set of $\{Q_{i},C_{i}\}$ between questions and best Cypher queries that contain the answers $\{A_{i}\}$. The best Cypher can be decided by reranking and for example filtering by recall and precision. We finetune an LLM $L_{1}$ with the given set of training data. The workflow is illustrated in Figure 2.
Figure 2: An example of creating ground-truth Cypher for a QA. In Step 1, few-shot LLM produces candidate entities which we ground with $\mathcal{G}$ in the DB with vector index. Step 2 shows part of the subgraph around the entity and answer nodes. With the DB, we execute the all one-hop, two-hop around each entity, and all length-two paths connecting the two entities in Step 3. We aggregate the hits and number of nodes for each query and rank them.
4.2 Grounded constrained decoding
Our $L_{1}$ has learned to generate graph pattern matching Cypher queries. However, there is still no guarantee that the generated Cypher is executable at inference time. The major bottlenecks are 1) syntactical correctness and 2) semantic correctness (i.e any type filters must correspond to existing node labels or edge types of $\mathcal{G}$). A common approach of generate-and-validate is both costly and slow and requires a sophisticated post-validation correction scheme to guarantee eventual correctness.
We use a new method of next-token constrained decoding at the level of logits processor at inference time. The constraints are applied at the token-level instead of word (or query keyword) level since the tokenizers used by the majorioty of modern LLMs do not have a one-to-one mapping from words and query keywords to the tokens.
Given a question at inference time, we first create a set of possible m-hop queries involving identified entities. This step is efficient since the schema of the graph is available. Let there be $\mathcal{Q}=\{\mathcal{Q}^{1},\cdots,\mathcal{Q}^{M}\}$ valid queries, each has tokenization $\mathcal{Q}^{k}=(n_{0},\cdots,n_{Q_{k}})\in\mathcal{V}^{Q_{k}}$ of variable length. When $L_{1}$ has generated $i$ tokens $(t_{0},\cdots,t_{i})$ and is generating a vector of logits $(l_{0},\cdots,l_{\mathcal{V}})$, our logits processor masks all invalid token with the value -$\infty$ by comparing with the $i+1^{\text{th}}$ token of all possible tokenized queries that match the initial $i$ tokens. The masking is defined as Equation 2 where $M_{i+1}$ is the set of valid tokens at $i+1^{\text{th}}$ position grounded in $\mathcal{G}$ via $\mathcal{Q}$.
$\tilde{\ell}_{i+1}^{(x)}=\begin{cases}\ell_{i+1}^{(x)},&\text{if }x\in M_{i+1}\\ -\infty,&\text{if }x\notin M_{i+1}\end{cases}\text{where }M_{i+1}=\{t\in\mathcal{V}|\exists\mathcal{Q}^{k}\in\mathcal{Q}\text{ , }\mathcal{Q}^{k}_{0:i+1}=(t_{0},\cdots,t_{k},t)\}$ (2)
The tokenizer for $L_{1}$ then applies decoding after softmax. By construction, our decoded query is executable. This constraint decoding is faithful and non-invasive for both greedy decoding and beam search with sufficiently large beams, as stated in Lemma 1. The formal statement and proof are provided in Appendix A. Unlike existing constrained generation methods[beurerkellner2024guidingllmsrightway, geng2024grammarconstraineddecodingstructurednlp], our approach does not require any formal grammar to be defined and is context-aware with respect to $\mathcal{G}$. An example of grounded constrained decoding is illustrated in Figure 3. Since our method is applied entirely on the logits before softmax, different final decoding such as greedy and beam search can be applied.
Lemma 1.
(Informal) If a query is invalid, it will not be generated. When using beam width = 1, constrained decoding acts as greedy decoding among valid queries. When beam width = M, exactly all valid queries are generated.
Figure 3: An example of grounded constrained decoding. For the given question, we tokenize all possible queries around it’s identified entities. At each step during generation, our logits processor masks out invalid tokens. For example, after "ichthyosis", the LLM would have generated the symbols ") which has the highest logit. Our processor masks it out since this predicate name: "X-linked ichthyosis" is invalid.
4.3 Finetuning LLM as local subgraph reasoner
Given any question for KG $\mathcal{G}$, our trained LLM $L_{1}$ produces guaranteed executable and grounded Cypher queries. If we perform beam search with width $m$, we are able to obtain $m$ valid queries with highest total probabilities. For the simplest questions and sparse $\mathcal{G}$, the subgraph produced by the best query may be exactly the answer nodes. However, for more difficult multi-hop questions on dense $\mathcal{G}$, the subgraph still contains other relevant nodes (and edges) that are not the answers. In order to obtain only the answer nodes by reasoning over the small subgraph, which often requires reasoning on the textual properties beyond graph patterns, we train an LLM $L_{2}$ to perform the task. We construct the subgraph by executing $m$ queries until some threshold on the size of the graph or tokenization required to encode the graph. The output of $L_{2}$ can be viewed as selecting or reranking the nodes in the input textualised graph. An example prompt we use is provided in Figure 4.
<|start_header_id|>user<|end_header_id|>
Given the information below, return the correct nodes for the following question:
What drugs target the CYP3A4 enzyme and are used to treat strongyloidiasis?
Retrieved information:
pattern: [’(x1:GeneOrProtein {name: "CYP3A4"})-[r1:ENZYME]-
(x2:Drug {name: "Ivermectin"})-[r2:INDICATION]-(x3:Disease {name: "strongyloidiasis"})’,
’(x1:Disease {name: "strongyloidiasis"})-[r1:INDICATION]-(x2:Drug {name: "Ivermectin"})’,
’(x1:GeneOrProtein {name: "CYP3A4"})-[r1:ENZYME]-(x2:Drug {name: "Ivermectin"})’]
name: Ivermectin
details: {’description’: ’Ivermectin is a broad-spectrum anti-parasite medication.
It was first marketed under the name Stromectol and used against worms (except tapeworms),
but, in 2012, it was approved for the topical treatment of head lice
infestations in patients 6 months of age and older,
and marketed under the name Sklice as well. ...
pattern: [’(x1:Disease {name: "strongyloidiasis"})-[r1:INDICATION]-
(x2:Drug {name: "Thiabendazole"})’]
name: Thiabendazole
details: {’description’: ’2-Substituted benzimidazole first introduced in 1962.
It is active against a variety of nematodes and is the drug of choice for
strongyloidiasis. It has CNS side effects and hepatototoxic potential.  ...
<|start_header_id|>model<|end_header_id|>
Figure 4: An example prompt that describes a local subgraph retrieved by Cypher queries around identified entities. This prompt contains both textual information and patterns used to retrieved them, which encodes topology information.
5 Experiments
Setup
We use Neo4j as the graph database. The default database configuration is used. For the main result, we using OpenAI text-embedding-ada-002[openai] as the text embedder $LM$ in Equation 1, OpenAI gpt-4o-mini as LLM $L_{0}$ for few-shot entity resolution., gemma2-9b-text2cypher[ozsoy2024text2cypher, team2024gemma] as our LLM $L_{1}$, Llama-3.1-8B-Instruct[grattafiori2024llama] as our LLM $L_{2}$. All experiments are run on a single 40GB A100 GPU. Additional detailed experiment setup is provided in Appendix B. All of our code is available on Github [^1] https://github.com/AlfredClemedtson/graphraft.
Datasets
We benchmark our method on the STaRK-prime and STaRK-mag datasets[wu2025stark] (license: CC-BY-4.0). stark-prime is a set of Q&As over a large biomedical knowledge graph PrimeKG[chandak2023building]. The questions mimic roles such as doctors, medical scientists and patients. It contains 10 node types and 18 edge types with rich textual properties on the nodes. With 129k nodes and 8 million edges and at the same time a high density (average node degree 125), it serves as a challenging and highly suitable dataset to benchmark retrieval and reasoning on large real-world KGs. stark-mag is a set of Q&As over ogbn-mag[hu2020open] that models a large academic citation network with relations between papers, authors, subject areas and institutions.
There is a large collection of Q&A datasets on graphs, we benchmark on datasets most suitable to the GraphRAG domain[bechler2025position] and elaborate on why we don’t benchmark on the others. WebQ[Berant2013SemanticPO], WebQSP[Yih2016TheVO], CWQ[talmor-berant-2018-web] ad GrailQA[gu2021beyond] are popular KBQA benchmarks containing SPARQL-answerable few-hop questions over Freebase[freebase], a database of general knowledge with URIs on nodes. Our problem setting has no requirement on the form of the questions and targets private KGs instead of such largely wikipedia-based KG which LLMs are explicitly trained on. Freebase has been deprecated since 2015. HotpotQA[yang2018hotpotqa] and BeerQA[qi2021answering] are few-hop questions over Wikidata[wikidata]. STaRK-amazon[wu2025stark] models properties of products (such as color) as nodes (with has-color relation) and the product co-purchasing graph itself is homogeneous.
5.1 Main results
We show our result in Table 1. The baselines range from pure vector-based retrievers to GraphRAG solutions to agentic methods. We use the four metrics originally proposed in [wu2025stark]. Hit@1 measure the ability of exact answering a right answer while the other three metrics provides more holistic view on the answer quality.
As is shown in Table 1, GraphRAFT gives best results on all metrics on both STaRK-prime and STaRK-mag. Even without using $L_{2}$ to reason over the local subgraph, our $L_{1}$ when used to retrieve nodes using generated Cypher queries (up to 20 nodes, for measuring recall), already gives better metrics than all SOTA methods on STaRK-prime.
Table 1: Main table of our results against previous baselines. Bold and underline represent the best method, underline represents the second best.
STARK-PRIME STARK-MAG
Hit@1 Hit@5 R@20 MRR Hit@1 Hit@5 R@20 MRR
BM25[Robertson2009ThePR] 12.75 27.92 31.25 19.84 25.85 45.25 45.69 34.91
voyage-l2-instruct[voyageai] 10.85 30.23 37.83 19.99 30.06 50.58 50.49 39.66
GritLM-7b[muennighoff2024generative] 15.57 33.42 39.09 24.11 37.90 56.74 46.40 47.25
multi-ada-002[openai] 15.10 33.56 38.05 23.49 25.92 50.43 50.80 36.94
ColBERTv2[santhanam-etal-2022-colbertv2] 11.75 23.85 25.04 17.39 31.18 46.42 43.94 38.39
Claude3 Reranker (10%) 17.79 36.90 35.57 26.27 36.54 53.17 48.36 44.15
GPT4 Reranker (10 %) 18.28 37.28 34.05 26.55 40.90 58.18 48.60 49.00
HybGRAG[lee2024hybgraghybridretrievalaugmentedgeneration] 28.56 41.38 43.58 34.49 65.40 75.31 65.70 69.80
AvaTaR[wu2025avatar] 18.44 36.73 39.31 26.73 44.36 59.66 50.63 51.15
MoR[lei2025mixturestructuralandtextualretrievaltextrich] 36.41 60.01 63.48 46.92 58.19 78.34 75.01 67.14
KAR[xia2025knowledgeawarequeryexpansionlarge] 30.35 49.30 50.81 39.22 50.47 65.37 60.28 57.51
MFAR[li2024multifieldadaptiveretrieval] 40.9 62.8 68.3 51.2 49.00 69.60 71.79 58.20
GraphRAFT w/o $L_{2}$ 52.12 71.55 75.52 60.72 62.63 86.68 88.88 73.03
GraphRAFT 63.71 75.39 76.39 68.99 69.64 84.32 89.12 76.24
5.2 Impact of constrained decoding and scaling with training data
We measure the benefit of applying grounded constrained decoding to using our model without it. We also examine how well the method scales with the availability of training data with and without constrained decoding. We measure the metrics directly on the output Cypher queries, executing the ones with highest probabilities first, until there are 20 nodes. To accurately evaluate LLM $L_{1}$ and constrained decoding we do not apply LLM $L_{2}$.
As can be seen in Table 2, when constrained decoding is not used and the model is trained on 100% of available data, it gives slightly lower metrics. When we only use 10% of the training data, our method only shows a slight decrease in all metrics. However, when used without constrained decoding, the drop becomes larger (e.g 16% for Recall@20). This suggests that our method is both extremely sample efficient and scales well with more training Q&As. The advantage of constrained decoding is the most significant when training data is scarce, which is common in any real-world setting. The final row uses the gemma2-9b-text2cypher model without any finetuning on STaRK-prime. It is not able to out-of-the-box answer any question and applying constrained decoding without finetuning the LLM at all already gives us results close to several baselines.
Table 2: Metrics on STaRK-prime using 10% of validation data. Percentage of train data used is specified next to the method. Numbers in brackets representing using the method without applying constrained decoding. No schema is provided in prompt and response executed as is in all queries.
Method (% Training data used) Hit@1 Hit@5 Recall@20 MRR
Finetuned, 100% 44.20(43.75) 69.20(63.39) 75.87(69.83) 0.5528(0.5255)
Finetuned, 10% 41.07(35.71) 66.07(54.91) 76.11(60.53) 0.5197(0.4400)
LLM, 0% 14.73(0.0) 23.21(0.0) 27.53(0.0) 0.1865(0.0)
5.3 The use of query engine
The use of query engines to optimise query plans on DBs has always been one of the main advantages of DB systems. Table 3 shows an example executed query plan that is optimal according to the query planner. It first fetches nodes of the type Drug from the node label indexes and then traverses along the ENZYME relationship type. It then filter the joined records with predicates on x1. Afterwards it performs the similar traversal from x2 to x3 along INDICATION and filter on x3. Intuitively, the optimal plan finds the all x2:Drug nodes and filter down twice by joining the two ends.
An alternative valid but suboptimal query plan is provided in Appendix C Table 4. It starts with scanning all x3:Disease nodes and traverses towards x2 and then x1. It happens to be suboptimal due to the exploding neighbourhood of Disease nodes along the INDICATION relationship type. The query engine with it’s cardinality estimator therefore rules out executing this sequence of operators.
This analysis serves as an example of confirming the advantage of using a query engine. As we have pointed out, any of the existing GraphRAG work that iteratively traverses the graph step-by-step is not able to leverage the query engine since the order of execution is fixed. Therefore, extremely inefficient retrieval (such as the ordering shown in Table 4) is possible and unavoidable.
Table 3: The optimal execution plan for an example retrieval query: MATCH (x1:GeneOrProtein {name: "CYP3A4"})-[r: ENZYME]-(x2: Drug)-[r2: INDICATION]-(x3: Disease {name: "strongyloidiasis"}) RETURN x2.name. Operator are executed from the bottom up. Details represent the exact execution parameters. Estimated Rows represent expected rows produced. Rows represent actual rows produced. DB Hits measure the amount of work by the storage engine. Total database access is 103989, total allocated memory is 328 bytes.
Operator Id Details Estimated Rows Rows DB Hits
+ProduceResults 0 n‘x2.name‘ 63 4 0
+Projection 1 x2.name AS ‘x2.name‘ 63 4 4
|+Filter 2 x3.name = \$autostring_1 AND x3:Disease 63 4 34872
|+Expand(All) 3 (x2)-[r2:INDICATION]-(x3) 1255 17432 17432
|+Filter 4 x1.name = \$autostring_0 AND x1:GeneOrProtein 532 1932 25132
|+Expand(All) 5 (x2)-[r:ENZYME]-(x1) 10634 10634 18591
+NodeByLabelScan 6 x2:Drug 7957 7957 7958
5.4 Ablation study on choices of LLMs
We first verify that using an LLM $L_{0}$ for entity resolution is needed better than simpler methods such as k-nearest-neighbour (kNN) using text embeddings. We measure that by looking at the quality of the best Cypher created from entities identified by an LLM, 2NN and 5NN. Plots showing the quality of entity resolution by various criteria are shown in Appendix D Figure 5.
We also study the different choices of base LLMs for $L_{1}$. In addition to using a general gemma-based text2cypher model, we also finetune a google/gemma2-9b-it and a meta-llama/Llama-3.1-8B-Instruct model. Metrics on STaRK-prime suggest using a already-fintuned general text2cypher model offers some advantage, and gemma2 performs slightly better tham Llama-3.1. Finetuning any of these base models for $L_{1}$ already achieves beyond SOTA results even without $L_{2}$. The table is shown in Appendix D Table 5.
6 Conclusion
In this work, we introduce GraphRAFT, a simple and modular method that leverages graph DBs by retrieving from it using learnt provably correct and optimal Cypher queries. Our experiments show that GraphRAFT consistently achieves beyond SOTA results using smaller LLMs that fit into a single GPU. Our framework can be applied off-the-shelf to any KG in any domain stored in graph DBs. The finetuning process is sample-efficient and scales with more training data.
GraphRAFT is illustrated on graph DBs supporting Cypher. Exactly the same approach can be used for any other graph query language. Our finetuning process requires existing Q&A set. Future work that addresses these limitations will improve the general applicability of the method.
Appendix A Constrained decoding
We formally state the informal Lemma 1 as three separate formal statements and prove them.
Recall notations in Section 4.2 where $\mathcal{Q}=\{\mathcal{Q}^{1},\cdots,\mathcal{Q}^{M}\}$ is the set of $M$ valid queries for a question, each has tokenization $\mathcal{Q}^{k}=(n_{0},\cdots,n_{Q_{k}})\in\mathcal{V}^{Q_{k}}$ of variable length. We use $l$ and $p$ to denote regular logits and probabilities and $\tilde{l}$ and $\tilde{p}$ to denote those produced by constrained decoding.
Lemma 1.1.
If $Q\notin\mathcal{Q}$ and $Q$ has tokenization $(t_{0},\dots,t_{q})$, then $\tilde{p}(Q)=\tilde{p}(t_{q}|t_{q-1},\dots,t_{0})\times\dots\times\tilde{p}(t_{0})=0$.
Proof.
If $t_{0}$ is not a valid first token, $\mathcal{Q}^{k}_{0}\neq t_{0}$ for all $k\in[1,M]$. Let the logits for the first token produced by LLM $L_{1}$ be $(l_{0},\dots,l_{\mathcal{V}})$. Let the logits after applying constrained decoding be $(\tilde{l}_{0},\dots,\tilde{l}_{\mathcal{V}})$. Then $\tilde{l}_{n_{0}}=$-$\infty$ and the probability of generating token $t_{0}$ is $\tilde{p}(n_{0})=0$.
Now let $(t_{0},\dots,t_{i})$ be the longest subsequence of valid tokens for $Q$. This means there are valid queries $Q^{k_{0}},\dots,Q^{k_{n}}$ such that $Q^{k_{j}}_{0,i}=(t_{0},\dots,t_{i})$ for $j\in[0,n]$. Since this is the longest subsequence, $Q^{k_{j}}_{i+1}\neq t_{i+1}$. For any $j$, let the logits of $i+1^{\text{th}}$ token be $(l_{0},\dots,l_{\mathcal{V}})$ and $(\tilde{l}_{0},\dots,\tilde{l}_{\mathcal{V}})$ after constrained decoding. $\tilde{l}_{t_{i+1}}=$-$\infty$. Hence $\tilde{p}(t_{i+1}|t_{i},\dots,t_{0})=0$. Therefore, $\tilde{p}(Q)=\tilde{p}(t_{q}|t_{q-1},\dots,t_{0})\times\dots\times\tilde{p}(n_{0})=0$ in all cases. ∎
Lemma 1.2.
Suppose that LLM $L_{1}$ has generated a sequence of valid tokens $(t_{0},\dots,t_{k})$, logits for the next token $k+1$ are $(l_{0},\dots,l_{\mathcal{V}})$. If $t_{k+1}$ and $t^{\prime}_{k+1}$ are two valid next tokens and $l_{t_{k+1}}>l_{t^{\prime}_{k+1}}$, then $\tilde{p}(t_{k+1}|t_{k},\dots,t_{0})>\tilde{p}(t^{\prime}_{k+1}|t_{k},\dots,t_{0})$.
Proof.
Since the two next tokens are valid, $l_{t_{k+1}}=\tilde{l}_{t_{k+1}}$ and $l_{t^{\prime}_{k+1}}=\tilde{l}_{t^{\prime}_{k+1}}$. $\tilde{p}(t_{k+1}|t_{k},\dots,t_{0})=\text{softmax}(\tilde{l}_{0},\dots,\tilde{l}_{\mathcal{V}})_{t_{k+1}}=e^{\tilde{l}_{t_{k+1}}}/\sum_{i=0}^{\mathcal{V}}e^{\tilde{l}_{i}}>e^{\tilde{l}_{t^{\prime}_{k+1}}}/\sum_{i=0}^{\mathcal{V}}e^{\tilde{l}_{i}}=\tilde{p}(t^{\prime}_{k+1}|t_{k},\dots,t_{0})$ simply by the monotonicity of the exponential function. ∎
Lemma 1.3.
When beam width = M, LLM $L_{1}$ with constrained decoding generates queries $\{Q^{1},\dots,Q^{M}\}=\{\mathcal{Q}^{1},\dots,\mathcal{Q^{M}}\}$.
Proof.
By the contrapositive of Lemma 1.1, $\tilde{p}(Q)\neq 0$ implies $Q\in\mathcal{Q}$. Hence, any query generated under constrained decoding is a valid query. $\{Q^{1},\dots,Q^{M}\}\subseteq\{\mathcal{Q}^{1},\dots,\mathcal{Q^{M}}\}$. Equality follows from cardinality.
Appendix B Experimental details
All experiments are done on a single machine with one 40GB A100 GPU, 85G RAM 12vCPUs. We use Neo4j Community Edition database with default database configurations. Total time for reproducing results for one datasets (e.g STaRK-prime), from loading raw data into the database, finetuning and evaluting metrics, is around 18 hours.
B.1 Few-shot prompt for entity resolution
B.1.1 STaRK-prime
Question : "Which anatomical structures lack the expression of genes or proteins involved in the interaction with the fucose metabolism pathway?"
Answer : "fucose metabolism"
Question : "What liquid drugs target the A2M gene/protein and bind to the PDGFR-beta receptor?"
Answer : "A2M gene/protein|PDGFR-beta receptor"
Question : "Which genes or proteins are linked to melanoma and also interact with TNFSF8?"
Answer : "melanoma|TNFSF8"
LLM insturction: "You are a knowledgeable assistant which identifies medical entities in the given sentences. Separate entities using '|'."
B.1.2 STaRK-mag
Question : "Could you find research articles on the measurement of radioactive gallium isotopes disintegration rates?"
Answer : "FieldOfStudy:measurement of radioactive gallium isotopes disintegration rates"
Question : "What research on water absorption in different frequency ranges have been referenced or deemed significant in the paper entitled 'High-resolution terahertz atmospheric water vapor continuum measurements'"
Answer: "FieldOfStudy:water absorption in different frequency ranges
Paper:High-resolution terahertz atmospheric water vapor continuum measurements"
Question : "Publications by Point Park University authors on stellar populations in tidal tails"
[…]
Question : "Show me publications by A.J. Turvey on the topic of supersymmetry particle searches."
[…]
LLM instruction: "You are a smart assistant which identifies entities in a given questions. There are institutions, authors, fields of study and papers."
B.2 K-hop query path templates
For both STaRK-prime and STaRK-mag, we use three path templates which are 1-hop, 2-hop and length two paths that connect two entities to curate training Cypher queries. Figure 1 shows the query for 2-hop. For 1-hop we substitue the pattern matching with:
MATCH (src {name: srcName})-[r]-(tgt)
and for length-two paths connecting entities, our query is:
UNWIND \$src_names AS srcName1
UNWIND \$src_names AS srcName2
MATCH (src1 {name: srcName1})-[r1]-(tgt)-[r2]-(src2 {name: srcName2})
WHERE src1 <> src2
RETURN labels(src1)[0] AS label1, src1.name AS name1, type(r1) AS type1, labels(tgt)[0] AS label2, type(r2) AS type2, labels(src2)[0] AS label3, src2.name AS name3, count(DISTINCT tgt) AS totalCnt
B.3 Finetuning experimental setups
B.3.1 Training LLM $L_{1}$
For finetuning LLM $L_{1}$ we use huggingface transformer library with LoRA. For base model neo4j/text2cypher-gemma-2-9b-it-finetuned-2024v1 (license: apache-2.0) and google/gemma-2-9b-it (license: gemma), we have
[…]
end_of_generation_token = "<eos>"
bnb_config = BitsAndBytesConfig(load_in_4bit=True, bnb_4bit_use_double_quant=True, bnb_4bit_quant_type="nf4",
bnb_4bit_compute_dtype=torch.bfloat16, )
tokenizer.padding_side = "right"
model = AutoModelForCausalLM.from_pretrained(
model_dir,
quantization_config=bnb_config,
torch_dtype=torch.bfloat16,
attn_implementation="eager",
low_cpu_mem_usage=True,
lora_config = LoraConfig(r=64, lora_alpha=64, target_modules=None, lora_dropout=0.05, bias="none", task_type="CAUSAL_LM", )
For base model meta-llama/Llama-3.1-8B-Instruct (license: llama3.1) used in ablation, the differences are in the tokenization and padding:
[…]
end_of_generation_token = "<|eot_id|>"
tokenizer.pad_token = '<|finetune_right_pad_id|>'
tokenizer.padding_side = "right"
The supervised finetuning training loop is configured as:
sft_config = SFTConfig(auto_find_batch_size=True,
gradient_accumulation_steps=1, #8
dataset_num_proc=8,
num_train_epochs=1,
learning_rate=2e-5,
optim="paged_adamw_8bit",
max_seq_length=max_seq_len,
eval_strategy="epoch",
save_strategy="epoch",
logging_steps=10,
output_dir=model_save_dir,
load_best_model_at_end=True,
B.3.2 Training LLM $L_{2}$
For finetuning LLM $L_{2}$ we use unsloth[unsloth]. For the base model meta-llama/Llama-3.1-8B-Instruct (license: llama3.1), our configurations are:
[…]
EOS = "<|eot_id|>"
RIGHT_PAD_TOKEN = '<|finetune_right_pad_id|>'
ANSWER_SEPARATOR = '|'
PRIME_MAX_SEQUENCE_LENGTH = 15_000
MAG_MAX_SEQUENCE_LENGTH = 15_000
MAX_NEW_TOKENS = 100
[…]
model, tokenizer = FastLanguageModel.from_pretrained(
model_name=model_dir,
max_seq_length=max_sequence_length, # None
dtype=torch.bfloat16,
load_in_4bit=True,
tokenizer.padding = True
tokenizer.pad_token = RIGHT_PAD_TOKEN
tokenizer.padding_side = 'right'
model = FastLanguageModel.get_peft_model(
model=model,
r=64,
target_modules=["q_proj", "k_proj", "v_proj", "o_proj", "gate_proj", "up_proj", "down_proj", ],
lora_alpha=64,
bias="none",
sft_config = SFTConfig(
per_device_train_batch_size=1,
per_device_eval_batch_size=1,
# auto_find_batch_size=True,
dataset_num_proc=8,
bf16=True,
num_train_epochs=1,
gradient_accumulation_steps=4,
warmup_steps=5,
learning_rate=2e-5,
optim="adamw_8bit",
weight_decay=0.01,
lr_scheduler_type="linear",
logging_steps=1,
eval_strategy="epoch",
save_strategy="epoch",
output_dir=model_save_dir,
load_best_model_at_end=True,
Appendix C Analysis of alternative query plans
An alternative valid but suboptimal query plan for the same Cypher query in Table 3 is shown in Table 4. The query optimiser rules out executing this less optimal plan. Adhoc retrieval methods in GraphRAG that do not use the underlying databases either require essentially a rewrite of the query engine in-memory, or risks having suboptimal retrievals.
Table 4: A valid but suboptimal query plan for an example retrieval query: MATCH (x1:GeneOrProtein {name: "CYP3A4"})-[r: ENZYME]-(x2: Drug)-[r2: INDICATION]-(x3: Disease {name: "strongyloidiasis"}) RETURN x2.name. Compared with the optimal plan in Table 3, this plan has both more operators and more costly ones.
Operator Id Details Estimated Rows
+ProduceResults 0 n‘x2.name‘ 63
+Projection 1 x2.name AS ‘x2.name‘ 63
|+NodeHashJoin 2 x2 109441
||+Filter 3 x1.name = \$autostring_0 AND x1:GeneOrProtein 51795
||+Expand(All) 4 (x2)-[r2:ENZYME]-(x1) 19893
||+NodeByLabelScan 5 x2:Drug 7957
|+Filter 6 x2:Drug 7957
|+Expand(All) 7 (x3)-[r:INDICATION]-(x2) 52521
|+Filter 8 x3.name = \$autostring_1 51240
+NodeByLabelScan 9 x3:Disease 17080
Appendix D Ablation of choices of LLMs
Given $\{Q_{i},A_{i}\}$ as input data, we prepare a set of $\{Q_{i},C_{i}\}$ as a set of training data to finetune the LLM to generate optimal Cypher queries. We first few-shot prompt an LLM to identify entities and then reconcile the identified names against the DB. Next, by using the graph schema, we obtain the set of all possible Cypher queries with type predicates k-hop around or connecting the entity nodes. The quality of the identified entities is therefore implicitly measured by the best Cypher that it allows. Figure 5 shows that using an LLM for entity resolution does produce better entities that lead to more high-recall ground-truth Cypher queries with more accurate subgraphs returned.
((a)) The number of questions that map to good Cyphers when using different entity resolution methods.
Figure 5: The impact of entity resolution on the quality of Cypher queries.
We also finetune different base LLMs for $L_{1}$ as shown in Table 5.
Table 5: Metircs for STaRK-prime when using $L_{1}$ only. Two different base models are finetuned. Both finetuned models achieve beyond SOTA results on every metric.
STARK-PRIME
Hit@1 Hit@5 R@20 MRR
Llama3.1-8b-Instruct 43.88 64.44 70.60 52.85
Gemma2-9b-it 48.27 68.23 74.46 57.12